Problem acceleration of a package sliding at Point A is 3 m/s2. Assuming that the coefficient of package kinetic is the same for each section, determine the acceleration of the package at Point B. Hint: Use an angle Cartesian coordinate system
Answer #1

now we find the coefficient of the friction at package A

a=g(sin(theta)-\mucos(theta))

3=9.8(sin30-\mucos30)

0.31=0.5-\mu0.87

\mu0.87=0.2

coefficient of the friction =0.23

now we find the acceleration of the package B

acceleration a2=9.8(sin15-0.23*cos15)

=0.36 m/s^2